OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2).

OpenStudy (amistre64):

(0,-2) IS your intercept so all you need to do is determine the slope

OpenStudy (amistre64):

subtract your points and stack y/x for the slope

OpenStudy (anonymous):

Could you guide me through the first step please?

OpenStudy (amistre64):

(4 , 8) -(0,-2) ------ ? , ?

OpenStudy (anonymous):

(4, 10) ?

OpenStudy (amistre64):

good :) now slope = y/x; slope = 10/4, and reduce

OpenStudy (anonymous):

Would y = 5/2 x - 2 be the answer?

OpenStudy (amistre64):

now, since slope intercept forms is: y = slope x + intercept we just fill in the pieces

OpenStudy (amistre64):

thats it

OpenStudy (anonymous):

okay thanks :)

OpenStudy (amistre64):

good job

OpenStudy (anonymous):

So for this would I do Determine the slope of the line which passes through these two points: (-1, 3) and (6, 9). (-1 , 3) - (6, 9) ----- ? ?